The angle of elevation of a hot air balloon changes from 30

to 45 degrees in 2 mins. If the point of observation is 200m away from the take

off point. Determine the uniform rate of change of the angle of elevation in

rad/sec.


1
Expert's answer
2021-11-23T10:32:33-0500
"\omega=\frac{45-30}{2\pi(120)}=0.02\frac{rad}{s}"